OpenStudy (anonymous):

Sound. Sound intensity (loudness) varies inversely as the square of the distance from the source. If a rock band has a sound intensity of 100 decibels 30 feet away from the amplifier, find the sound intensity 60 feet away from the amplifier.

OpenStudy (anonymous):

Do you know the formula for sound?

OpenStudy (anonymous):

It's a variation problem. The formula they give me are: varies directly: d=kf varies inversely: I=k/R joint variation: g=khr^2 combined variation: y=kx/z

OpenStudy (anonymous):

okay, so we will be using l=k/r

OpenStudy (anonymous):

\[100=\frac{ k }{ 30 }\]

OpenStudy (anonymous):

3000=k

OpenStudy (anonymous):

thank you

OpenStudy (dan815):